# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 2.2 Potential kernel
$a(x,t):=\int_{0}^{t}\Big{(}{\mathbb{P}}(X^{0}_{s}=0)-{\mathbb{P}}(X^{x}_{s}=0)\Big{)}\mathrm{d}s\quad\mbox{and}\quad a(x):=\lim\limits_{t\to\infty}a(x,t).$ (2.14)
It is known that for a one-dimensional random walk the limit $a(x)\in[0,\infty)$ exists for all $x\in{\mathbb{Z}}$ (see [Spi01, T28.1]. We need a finer estimate that quantifies the convergence rate.
@@
(i) If $10<\alpha<$ 1, then for any $0<\kappa_{1}<\frac{1-\alpha}{\alpha}$ such that uniformly in $x$,
$a(x)-a(x,t)=O\big{(}t^{-\kappa_{1}}\big{)}.$ (2.15)
(ii)If $21\leq\alpha<$ 2, then there exist $\kappa_{1}>0$ and $\kappa_{2}<1$ such that
$a(x)-a(x,t)=O\big{(}t^{-\kappa_{1}}|x|^{\kappa_{2}}\big{)}.$ (2.16)
@@
$\int_{0}^{T}\frac{1-\cos(xu)}{u^{\kappa}}\mathrm{d}u=\left\{\begin{array}[]{ll}\displaystyle O\Big{(}T^{1-\kappa}\Big{)},&\text{if~{}}0<\kappa<1,\\[10.0pt] \displaystyle O\Big{(}\log(|x|T)\Big{)},&\text{if~{}}\kappa=1,\\[10.0pt] \displaystyle O\Big{(}|x|^{\theta+\kappa-1}T^{\theta}\Big{)},&\text{if~{}}1<\kappa<2.\end{array}\right.$ (2.17)
@@
$\int_{0}^{T}\frac{1-\cos(xu)}{u^{\kappa}}\mathrm{d}u=x^{\kappa-1}\int_{0}^{xT}\int_{0}^{u}\frac{\sin v}{u^{\kappa}}\mathrm{d}v\mathrm{d}u=x^{\kappa-1}\int_{0}^{xT}\int_{v}^{xT}\frac{\sin v}{u^{\kappa}}\mathrm{d}u\mathrm{d}v.$ (2.18)
@@
(2.18$)\leq\frac{x^{\kappa-1}}{1-\kappa}\int_{0}^{xT}(xT)^{1-\kappa}\sin v\mathrm{d}v\leq CT^{1-\kappa},\quad(0<\kappa<1).$
(2.18$)\leq\int_{0}^{xT}\sin v\log\frac{xT}{v}\mathrm{d}v\leq C\log(xT),\quad(\kappa=1).$
For $21<\kappa<$ 2, if $xT\geq 1$, then by the integrability of $\sin v/v^{\kappa-1}$ over $(0,\infty)$,
(2.18$)\leq\frac{x^{\kappa-1}}{\kappa-1}\int_{0}^{xT}\frac{\sin v}{v^{\kappa-1}}\mathrm{d}v\leq Cx^{\kappa-1}\leq Cx^{\kappa-1}(xT)^{\theta},\quad(\theta\geq 0).$
@@
(2.18$)\leq Cx^{\kappa-1}\int_{0}^{xT}v^{2-\kappa}\mathrm{d}v\leq Cx^{\kappa-1}(xT)^{3-\kappa}\leq Cx^{\kappa-1}(xT)^{\theta},\quad(\theta\leq 3-\kappa).$
@@
$\begin{array}[]{l}\displaystyle a(x)-a(x,t)=\frac{1}{2\pi}\int_{-\pi}^{\pi}\int_{t}^{\infty}\!\!\big{(}1-e^{ixu}\big{)}\phi_{s}(u)\mathrm{d}s\mathrm{d}u=\frac{1}{\pi}\!\int_{0}^{\pi}\!\frac{1-\cos(xu)}{1-\phi(u)}e^{-t\big{(}1-\phi(u)\big{)}}\mathrm{d}u.\end{array}$ (2.19)
@@
where $\varepsilon>0$ can be chosen arbitrary small. Since $t(1-\phi(u))\geq Ct(t^{-\frac{1-\varepsilon}{\alpha+\varepsilon}})^{\alpha+\varepsilon}=Ct^{\varepsilon}$ on $D_{2}$,
@@
For $10<\alpha<$ 1, choose $\varepsilon<1-\alpha$. Then, by Lemma 2.5
@@
For $21\leq\alpha<$ 2, choose $\varepsilon<\frac{2-\alpha}{2}$ and apply Lemma 2.5 with $\theta=2-\alpha-2\varepsilon$, yielding
@@
This is analogous to the relation between the Green’s function and the jump kernel of a transient random walk. We will use (2.20) for $x$=0. Observe that if the random walk is of rate $r$, then $\sum_{y}p(y)a(y)=r$.
@@
The results in this subsection extend to non-symmetric jump kernels. The only modification in the proofs is that the characteristic function is replaced by $\phi^{\text{ns}}$ in (2.9). In our arguments, the required estimates for a single random walk rely solely on the results from this subsection and the previous one (Remark 2.2). The remaining estimates, particularly those involving collision times, can be reduced to the first hitting times of difference process between two random walks. Notably, the jump kernel of the difference process is always symmetric. Consequently, the main result, Theorem 1.1, remains valid even for non-symmetric $p(\cdot)$.
@@ 2.3 First hitting time
[…]
Our goal is to derive asymptotic estimates for the tail probability ${\mathbb{P}}(\tau_{x}>t)$. A classic approach (which we leave the details to readers) is to apply the Laplace transform to the last-exit formula and then invoke a Tauberian theorem, which yields
${\mathbb{P}}(\tau_{x}>t)\sim\frac{a(x)}{c_{\alpha}g_{\alpha}(0)l(t)},$ (2.22)
where $c_{\alpha}=1$ for $0<\alpha\leq 1$ and $c_{\alpha}=\Gamma(1/\alpha)\Gamma(2-1/\alpha)$ for $21<\alpha<$ 2. We leave the details to readers. Nevertheless, to incorporate these estimates into our ODE framework, we require a higher order estimate of the tail probability. Fortunately, for $\alpha\leq 1$ the function ${\mathbb{P}}(\tau_{x}>t)$ is slowly varying, which allows us to obtain the refined result Proposition 2.11.
@@
$j(t):=\sum_{x\in{\mathbb{Z}}}p(x){\mathbb{P}}(\tau_{x}>t)=\sum_{x\neq 0}p(x){\mathbb{P}}(\tau_{x}>t),$ (2.23)
which represents that the random walk jumps away from 0 initially and does not hit 0 by time $t$. Recall the last-exit formulae that decompose at the last time the walk is at 00,
$\begin{array}[]{l}\displaystyle 1={\mathbb{P}}(X^{0}_{t}=0)+\int_{0}^{t}{\mathbb{P}}(X^{0}_{s}=0)j(t-s)\mathrm{d}s,\\[5.0pt] \displaystyle{\mathbb{P}}(\tau_{x}\leq t)={\mathbb{P}}(X^{x}_{t}=0)+\int_{0}^{t}{\mathbb{P}}(X^{x}_{s}=0)j(t-s)\mathrm{d}s,\quad(x\neq 0),\end{array}$ (2.24)
We begin by estimating $j(t)$. Recall that $g_{\alpha}$ is the density function of the limiting $\alpha$-stable distribution, and recall from (1.7) that $l(t)=\int_{0}^{t}\frac{1}{B_{s}}\mathrm{d}s$ and $\ell(t)=B_{t}\hskip 1.0ptl(t)/t$.
@@
For $j(t)$ defined in (2.23) and for $\alpha\leq 1$, one has
$\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t)=j(t)=\frac{1}{g_{\alpha}(0)l(t)}\Big{(}1+O(1/\ell(t))\Big{)}.$ (2.25)
Proof Since $j(t)=\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t)$ is decreasing in $t$, the last-exit formula (2.24) implies
@@
where in the last inequality, for the integral from $t$ to $2t$ we used $j(t)\leq\frac{1}{g_{\alpha}(0)l(t)}$ and then $\int_{0}^{t}\frac{1}{l(s)}\mathrm{d}s\sim\frac{t}{l(t)}$. Note that $B_{t}\gg\ell(t)$. The estimate (2.25) follows.
@@
For $j(t)$ defined in (2.23) and for $\alpha\leq 1$, there exists $C$ such that for $s<t$ and any $\varepsilon>0$,
$j(s)-j(t)\leq\frac{C}{l(t)\ell(t)}(t/s)^{\frac{1}{\alpha}-1+\varepsilon}.$ (2.26)
@@
$l(t)-l(s)=\frac{t^{1/\alpha}}{B_{t}}\int_{s}^{t}\frac{u^{1/\alpha}/B_{u}}{t^{1/\alpha}/B_{t}}\frac{1}{u^{1/\alpha}}\mathrm{d}u\leq\frac{t^{1/\alpha}}{B_{t}}\int_{s}^{t}\frac{(t/u)^{\varepsilon}}{u^{1/\alpha}}\mathrm{d}u\leq\frac{t\hskip 1.0pt\hskip 1.0pt(t/s)^{1/\alpha-1+\varepsilon}}{B_{t}}.$ (2.27)
@@
Inserting these bounds into (2.25) yields that for any $\varepsilon>0$,
@@
$\Delta(x,t):={\mathbb{P}}(X_{t}^{0}=0)-{\mathbb{P}}(X_{t}^{x}=0)$ (2.28)
so that the last-exit formula (2.24) yields, for $x\neq 0$,
${\mathbb{P}}(\tau_{x}>t)=\Delta(x,t)+\int_{0}^{t}\Delta(x,s)j(t-s)\mathrm{d}s.$ (2.29)
A uniform bound on the transition probability, combined with Proposition 2.1, shows that
$0\leq\Delta(x,t)\leq CB_{t}^{-1}(1\wedge|x|B_{t}^{-1}).$ (2.30)
@@
Let $(X^{x}_{t})_{t\geq 0}$ be the heavy-tailed random walk with increments in the domain of attraction of an $\alpha$-stable law for some $0<\alpha\leq 1$. Recall the first hitting time $\tau_{x}$ from (2.21). Recall $B_{t}$ from (1.5), and recall $l(t)$ and $\ell(t)$ from (1.7). Then, for $x\neq 0$ and any $\varepsilon>0$, there exist some $\kappa>0$ such that
${\mathbb{P}}(\tau_{x}>t)=\left\{\begin{array}[]{ll}\displaystyle\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big{(}\frac{a(x)}{l(t)\ell(t)}+\frac{|x|}{t^{\kappa}}\Big{)},&\alpha=1;\\[12.0pt] \displaystyle\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big{(}\frac{a(x)}{l(t)\ell(t)}+\frac{t^{\varepsilon}}{\ell(t)}\Big{)},&0<\alpha<1.\end{array}\right.$ (2.31)
Proof Rewrite (2.29) as
${\mathbb{P}}(\tau_{x}>t)=\int_{0}^{t}\Delta(x,s)j(t)\mathrm{d}s+\int_{0}^{t}\Delta(x,s)\big{(}j(t-s)-j(t)\big{)}\mathrm{d}s+\Delta(x,t).$ (2.32)
@@
$\int_{0}^{t}\Delta(x,s)j(t)\mathrm{d}s=a(x)j(t)+\big{(}a(x,t)-a(x)\big{)}j(t)=\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big{(}\frac{a(x)}{l(t)\ell(t)}+\frac{|x|^{\kappa_{2}}}{t^{\kappa_{1}}l(t)}\Big{)},$
where $\kappa_{1}>0$ and $\kappa_{2}\in(0,1)$ if $\alpha=1$, and $\kappa_{1}$ is any constant in $(0,\frac{1-\alpha}{\alpha})$ and $\kappa_{2}=0$ if $10<\alpha<$ 1. Secondly, applying Lemma 2.10 to $j(t-s)-j(t)$ for $s\in(0,t/2)$,
$\int_{0}^{\frac{t}{2}}\Delta(x,s)\big{(}j(t-s)-j(t)\big{)}\mathrm{d}s\leq\frac{C}{l(t)\ell(t)}\int_{0}^{\infty}\Delta(x,s)\mathrm{d}s=\frac{Ca(x)}{l(t)\ell(t)}.$
For $s\in(t/2/,t)$, Karamata’s theorem implies $\int_{0}^{t}j(s)\mathrm{d}s\asymp t/l(t)$, and thus by (2.30)
$\int^{t}_{\frac{t}{2}}\Delta(x,s)\big{(}j(t-s)-j(t)\big{)}\mathrm{d}s\leq\frac{C}{B_{t}}\Big{(}\frac{|x|}{B_{t}}\wedge 1\Big{)}\int_{0}^{\frac{t}{2}}\big{(}j(s)+j(t)\big{)}\mathrm{d}s\leq\frac{C}{\ell(t)}\Big{(}\frac{|x|}{B_{t}}\wedge 1\Big{)}.$
For $\alpha=1$, since both $l$ and $\ell$ are SVFs, there exists $\kappa\in(0,1)$ such that both $\frac{|x|^{\kappa_{2}}}{t^{\kappa_{1}}l(t)}$ and $\frac{|x|}{\ell(t)B_{t}}$ are bounded by $|x|/t^{\kappa}$. For $10<\alpha<$ 1 and for any $\varepsilon>0$ small enough, choose $\kappa_{1}=1/\alpha-1-2\varepsilon$. Since now $l$ is a SVF and $\ell$ is $(1/\alpha\!-\!1)$-regularly varying, $\frac{1}{t^{\kappa_{1}}l(t)}$ is bounded by $t^{\varepsilon}/\ell(t)$. Inserting all these bounds into (2.32), we obtain (2.31).
Since $\ell(t)=B_{t}\hskip 1.0ptl(t)/t$, we observe that $B_{t}\gg\ell(t)^{\theta}$ if $0<\theta<\frac{1}{1-\alpha}$ ( (with the convention that if $\alpha=1$ then $\frac{1}{1-\alpha}=\infty$). The following lemma gives a tail bound on the hitting time when the starting position is relatively far from the origin.
@@
${\mathbb{P}}(\tau_{x}\leq t)\leq C\hskip 1.0pt\ell(t)^{-1+(1-\alpha)\theta+\varepsilon}.$ (2.33)
Proof We use the last-exit formula (2.24). Recall that $L(B_{t})/B_{t}^{\alpha}\sim 1/t$. By the Potter bound
@@
Set $s:=t/\ell(t)^{\alpha\theta}$. Since the polynomial power in $\ell(t)^{\alpha\theta}$ is $\frac{1-\alpha}{\alpha}\cdot\alpha\theta<1$, we have $s\gg 1$. We now split the integration in (2.24) into three regions. For $u\in(0,s]$, by the local large deviation (2.10), Karamata’s theorem, and $j(t-u)\leq C/l(t)$,
$\int_{0}^{s}P(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\int_{0}^{s}\frac{CuL\big{(}B_{t}/\ell(t)^{\theta}\big{)}}{B_{u}B_{t}^{\alpha}/\ell(t)^{\alpha\theta}}\cdot\frac{1}{l(t)}\mathrm{d}u\leq\frac{Cs^{2}\ell(t)^{\alpha\theta+\varepsilon}}{tB_{s}\hskip 1.0ptl(t)}\leq\frac{C\hskip 1.0pt(t/s)^{\frac{1}{\alpha}-2+\varepsilon}}{\ell(t)^{1-\alpha\theta-\varepsilon}}.$
For $u\in(s,t-s]$, we bound $j(t-u)\leq j(s)\leq C/l(s)$, and use (2.27) to obtain
$\int^{t-s}_{s}P(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\frac{C}{l(s)}\big{(}l(t)-l(s)\big{)}\leq\frac{C}{\ell(t)}(t/s)^{\frac{1}{\alpha}-1+\varepsilon}.$
@@
$\int_{t-s}^{t}P(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\frac{C}{B_{t}}\int_{0}^{s}j(u)\mathrm{d}u\leq\frac{Cs}{B_{t}l(s)}\leq\frac{C}{\ell(t)}(t/s)^{-1+\varepsilon}.$
Combining the three bounds above and inserting $t/s=\ell(t)^{\alpha\theta}$ gives

### v2

@@ 2.2 Potential kernel
$a(x,t):=\int_{0}^{t}\Big({\mathbb{P}}(X^{0}_{s}=0)-{\mathbb{P}}(X^{x}_{s}=0)\Big)\mathrm{d}s\quad\mbox{and}\quad a(x):=\lim\limits_{t\to\infty}a(x,t).$ (33)
It is known that for a one-dimensional random walk the limit $a(x)\in[0,\infty)$ exists for all $x\in{\mathbb{Z}}$ (see (Spi, 01, T28.1). We need a finer estimate that quantifies the convergence rate.
@@
(i) If $0<\alpha<1$, then for any $0<\kappa_{1}<\frac{1-\alpha}{\alpha}$, we have uniformly in $x$,
$a(x)-a(x,t)=O\big(t^{-\kappa_{1}}\big).$ (34)
(ii) If $1\leq\alpha<2$, then there exist $\kappa_{1}>0$ and $\kappa_{2}\in(\alpha/2,1)$ such that
$a(x)-a(x,t)=O\big(t^{-\kappa_{1}}|x|^{\kappa_{2}}\big).$ (35)
@@
$\int_{0}^{T}\frac{1-\cos(xu)}{u^{\kappa}}\mathrm{d}u=\left\{\begin{array}[]{ll}\displaystyle O\Big(T^{1-\kappa}\Big),&\text{if~}0<\kappa<1,\\[10.0pt] \displaystyle O\Big(\log(|x|T)\vee 1\Big),&\text{if~}\kappa=1,\\[10.0pt] \displaystyle O\Big(|x|^{\theta+\kappa-1}T^{\theta}\Big),&\text{if~}1<\kappa<2.\end{array}\right$. (36)
@@
$\int_{0}^{T}\frac{1-\cos(xu)}{u^{\kappa}}\mathrm{d}u=x^{\kappa-1}\int_{0}^{xT}\int_{0}^{u}\frac{\sin v}{u^{\kappa}}\mathrm{d}v\mathrm{d}u=x^{\kappa-1}\int_{0}^{xT}\int_{v}^{xT}\frac{\sin v}{u^{\kappa}}\mathrm{d}u\mathrm{d}v.$ (37)
@@
[…]
For $1<\kappa<2$, if $xT\geq 1$, then by the integrability of $\sin v/v^{\kappa-1}$ over $(0,\infty)$,
[…]
@@
[…]
@@
$\begin{array}[]{l}\displaystyle a(x)-a(x,t)=\frac{1}{2\pi}\int_{-\pi}^{\pi}\int_{t}^{\infty}\!\!\big(1-e^{ixu}\big)\phi_{s}(u)\mathrm{d}s\mathrm{d}u=\frac{1}{\pi}\!\int_{0}^{\pi}\!\frac{1-\cos(xu)}{1-\phi(u)}e^{-t\big(1-\phi(u)\big)}\mathrm{d}u.\end{array}$ (38)
@@
Since $t(1-\phi(u))\geq Ct(t^{-\frac{1-\varepsilon}{\alpha+\varepsilon}})^{\alpha+\varepsilon}=Ct^{\varepsilon}$ on $D_{2}$,
@@
For $0<\alpha<1$, choose $\varepsilon<1-\alpha$. Then, by Lemma 2.5
@@
For $1\leq\alpha<2$, choose $\varepsilon<\frac{2-\alpha}{2}$ and apply Lemma 2.5 with $\theta=2-\alpha-2\varepsilon$, yielding
@@
This is analogous to the relation between the Green’s function and the jump kernel of a transient random walk. We will use (39) for xx=0. Observe that if the random walk is of rate $r$, then $\sum_{y}p(y)a(y)=r$.
@@
In the non-symmetric case, where the characteristic function is replaced by $\phi^{\mathrm{ns}}$ in (30), an equation analogous to (38) holds for the symmetrized quantity a(x)a(x,t)+a(x)a(x,t)a(x)-a(x,t)+a(-x)-a(-x,t). Hence, Proposition 2.4 extends to this symmetrized form. However, deriving the result for a(x)a(x,t)a(x)-a(x,t) alone presents certain technical difficulties that may require a more refined analysis, particularly in the case $\alpha=1$ when the mean does not exist; see, for instance, (Ber, 19, Section 4) for further discussion. We note that, in our arguments, the required estimates for a single random walk rely solely on the results from this subsection and the previous one (Remark 2.2). The remaining estimates, particularly those involving collision times, can be reduced to the first hitting times of difference process between two random walks. Notably, the jump kernel of the difference process is always symmetric. Consequently, the main result, Theorem 1.3, would continue to hold provided Proposition 2.4 can be extended to non-symmetric jump kernels $p(\cdot)$.
@@ 2.3 First hitting time
$\tau_{0,x}:=\tau_{x}:=\inf\{t:X_{t}^{x}=0\text{~and~}X^{x}_{s}\neq 0~\text{ for some~}0<s<t\}.$ (40)
Our goal is to derive asymptotic estimates for the tail probability ${\mathbb{P}}(\tau_{x}>t)$. A classic approach (which we leave the details to readers) is to apply the Laplace transform to the last-exit formula and then invoke a Tauberian theorem (BGT, 89, Theorem 1.7.1), which yields
${\mathbb{P}}(\tau_{x}>t)\sim\frac{a(x)}{c_{\alpha}g_{\alpha}(0)l(t)},$ (41)
where $c_{\alpha}=1$ for $0<\alpha\leq 1$ and $c_{\alpha}=\Gamma(1/\alpha)\Gamma(2-1/\alpha)$ for $1<\alpha<2$. We leave the details to readers. Nevertheless, to incorporate these estimates into our ODE framework, we require a higher order estimate of the tail probability. Fortunately, for $\alpha\leq 1$ the function ${\mathbb{P}}(\tau_{x}>t)$ is slowly varying, which allows us to obtain the refined result Proposition 2.11.
@@
$j(t):=\sum_{x\in{\mathbb{Z}}}p(x){\mathbb{P}}(\tau_{x}>t)=\sum_{x\neq 0}p(x){\mathbb{P}}(\tau_{x}>t),$ (42)
which represents that the random walk jumps away from 0 initially and does not hit 0 by time $t$. Recall the last-exit formulae that decompose at the last time the walk is at 0,
$\begin{array}[]{l}\displaystyle 1={\mathbb{P}}(X^{0}_{t}=0)+\int_{0}^{t}{\mathbb{P}}(X^{0}_{s}=0)j(t-s)\mathrm{d}s,\\[5.0pt] \displaystyle{\mathbb{P}}(\tau_{x}\leq t)={\mathbb{P}}(X^{x}_{t}=0)+\int_{0}^{t}{\mathbb{P}}(X^{x}_{s}=0)j(t-s)\mathrm{d}s,\quad(x\neq 0),\end{array}$ (43)
We begin by estimating $j(t)$. Recall that $g_{\alpha}$ is the density function of the limiting $\alpha$-stable distribution, and recall from (7) that $l(t)=\int_{0}^{t}\frac{1}{B_{s}}\mathrm{d}s$ and $\ell(t)=B_{t}\hskip 1.0ptl(t)/t$.
@@
For $j(t)$ defined in (42) and for $\alpha\leq 1$, one has
$\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t)=j(t)=\frac{1}{g_{\alpha}(0)l(t)}\Big(1+O(1/\ell(t))\Big).$ (44)
Proof Since $j(t)=\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t)$ is decreasing in $t$, the last-exit formula (43) implies
@@
where in the last inequality, for the integral from $t$ to $2t$ we used $j(t)\leq\frac{1}{g_{\alpha}(0)l(t)}$ and then $\int_{0}^{t}\frac{1}{l(s)}\mathrm{d}s\sim\frac{t}{l(t)}$. Note that $B_{t}\gg\ell(t)$. The estimate (44) follows.
@@
For $j(t)$ defined in (42) and for $\alpha\leq 1$, there exists $C$ such that for $s<t$ and any $\varepsilon>0$,
$j(s)-j(t)\leq\frac{C}{l(t)\ell(t)}\left(\frac{t}{s}\right)^{\frac{1}{\alpha}-1+\varepsilon}.$ (45)
@@
$\begin{array}[]{l}\displaystyle l(t)-l(s)=\frac{t^{1/\alpha}}{B_{t}}\int_{s}^{t}\frac{u^{1/\alpha}/B_{u}}{t^{1/\alpha}/B_{t}}\frac{1}{u^{1/\alpha}}\mathrm{d}u\leq\frac{t^{1/\alpha}}{B_{t}}\int_{s}^{t}\frac{(t/u)^{\varepsilon}}{u^{1/\alpha}}\mathrm{d}u\\[10.0pt] \hskip 142.26378pt\displaystyle\leq\frac{t}{B_{t}}\left(\frac{t}{s}\right)^{\frac{1}{\alpha}-1+\varepsilon}=\frac{l(t)}{\ell(t)}\left(\frac{t}{s}\right)^{\frac{1}{\alpha}-1+\varepsilon}.\end{array}$ (46)
@@
Inserting these bounds into (44) yields that for any $\varepsilon>0$,
@@
$\Delta(x,t):={\mathbb{P}}(X_{t}^{0}=0)-{\mathbb{P}}(X_{t}^{x}=0)\geq 0,$ (47)
so that the last-exit formula (43) yields, for $x\neq 0$,
${\mathbb{P}}(\tau_{x}>t)=\Delta(x,t)+\int_{0}^{t}\Delta(x,s)j(t-s)\mathrm{d}s.$ (48)
Note that a uniform bound on the transition probability ${\mathbb{P}}(X^{x}_{t}=0)\leq C/B_{t}$ implies the uniform bound $\Delta(x,t)\leq C/B_{t}$. This, combined with Proposition 2.1, shows that
$0\leq\Delta(x,t)\leq CB_{t}^{-1}(|x|B_{t}^{-1}\wedge 1).$ (49)
@@
Let $(X^{x}_{t})_{t\geq 0}$ be the heavy-tailed random walk with increments in the domain of attraction of an $\alpha$-stable law for some $0<\alpha\leq 1$. Recall the first hitting time $\tau_{x}$ from (40). Recall $B_{t}$ from (5), and recall $l(t)$ and $\ell(t)$ from (7). Then, for $x\neq 0$ and any $\varepsilon>0$, there exists some $\kappa>0$ such that
${\mathbb{P}}(\tau_{x}>t)=\left\{\begin{array}[]{ll}\displaystyle\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big(\frac{a(x)}{l(t)\ell(t)}+\frac{|x|}{t^{\kappa}}\Big),&\alpha=1;\\[12.0pt] \displaystyle\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big(\frac{a(x)}{l(t)\ell(t)}+\frac{t^{\varepsilon}}{\ell(t)}\Big),&0<\alpha<1.\end{array}\right$. (50)
Proof Rewrite (48) as
${\mathbb{P}}(\tau_{x}>t)=\int_{0}^{t}\Delta(x,s)j(t)\mathrm{d}s+\int_{0}^{t}\Delta(x,s)\big(j(t-s)-j(t)\big)\mathrm{d}s+\Delta(x,t).$ (51)
@@
$\int_{0}^{t}\Delta(x,s)j(t)\mathrm{d}s=a(x)j(t)+\big(a(x,t)-a(x)\big)j(t)=\frac{a(x)}{g_{\alpha}(0)l(t)}+O\Big(\frac{a(x)}{l(t)\ell(t)}+\frac{|x|^{\kappa_{2}}}{t^{\kappa_{1}}l(t)}\Big).$ (52)
In the equation above, if $\alpha=1$, then $\kappa_{1}>0,\kappa_{2}\in(0,1)$ and thus $\frac{|x|^{\kappa_{2}}}{t^{\kappa_{1}}l(t)}\leq C\frac{|x|}{t^{\kappa}}$ for $\kappa\in(0,\kappa_{1})$; and if $0<\alpha<1$, then $\kappa_{1}\in(0,\frac{1-\alpha}{\alpha}),\kappa_{2}=0$ and thus
where $\kappa_{1}>0$ and $\kappa_{2}\in(0,1)$ if $\alpha=1$, and $\kappa_{1}\in(0,\frac{1-\alpha}{\alpha})$ and $\kappa_{2}=0$ if $0<\alpha<1$. Secondly, applying Lemma 2.10 to j(ts)j(t)j(t-s)-j(t) for $s\in(0,t/2)$,
$\int_{0}^{\frac{t}{2}}\Delta(x,s)\big(j(t-s)-j(t)\big)\mathrm{d}s\leq\frac{C}{l(t)\ell(t)}\int_{0}^{\infty}\Delta(x,s)\mathrm{d}s=\frac{Ca(x)}{l(t)\ell(t)}.$ (53)
For $s\in(t/2/,t)$, Karamata’s theorem implies $\int_{0}^{t}j(s)\mathrm{d}s\asymp t/l(t)$, and thus by (49)
$\int^{t}_{\frac{t}{2}}\Delta(x,s)\big(j(t-s)-j(t)\big)\mathrm{d}s\leq\frac{C}{B_{t}}\Big(\frac{|x|}{B_{t}}\wedge 1\Big)\int_{0}^{\frac{t}{2}}\big(j(s)+j(t)\big)\mathrm{d}s\leq\frac{C}{\ell(t)}\Big(\frac{|x|}{B_{t}}\wedge 1\Big).$ (54)
Recall that $B_{t}\geq\ell(t)$ for large $t$. (49) implies that $\Delta(x,t)\leq C\ell(t)^{-1}(|x|B_{t}^{-1}\wedge 1)$.
For $\alpha=1$, since both $l$ and $\ell$ are SVFs, there exists $\kappa\in(0,1)$ such that both $\frac{|x|^{\kappa_{2}}}{t^{\kappa_{1}}l(t)}$ and $\frac{|x|}{\ell(t)B_{t}}$ are bounded by $|x|/t^{\kappa}$. Using this bound for the r.h.s of (52) and (54), and inserting (52)-(54) into (51), we obtain (50) for $\alpha=1$. For $0<\alpha<1$ and for any $\varepsilon>0$ small enough, recall $\kappa_{2}=0$ and choose $\kappa_{1}=1/\alpha-1-2\varepsilon$ in (52). Since now $l$ is a SVF and $\ell$ is $(1/\alpha\!-\!1)$-regularly varying, $\frac{1}{t^{\kappa_{1}}l(t)}$ is bounded by $t^{\varepsilon}/\ell(t)$. Thus, inserting (52)-(54) into (51), we obtain (50) for $0<\alpha<1$.
Since $\ell(t)=B_{t}\hskip 1.0ptl(t)/t$, we observe that $B_{t}\gg\ell(t)^{\theta}$ if $0<\theta<\frac{1}{1-\alpha}$ (with the convention that $\frac{1}{1-\alpha}=\infty$ if $\alpha=1$). The following lemma gives a tail bound on the hitting time when the starting position is relatively far from the origin.
@@
${\mathbb{P}}(\tau_{x}\leq t)\leq C\hskip 1.0pt\ell(t)^{-1+(1-\alpha)\theta+\varepsilon}.$ (55)
Proof We use the last-exit formula (43). Recall from (6) that $L(B_{t})/B_{t}^{\alpha}\sim 1/t$. By the Potter bound
@@
Set $s:=t/\ell(t)^{\alpha\theta}$. Since the polynomial power in $\ell(t)^{\alpha\theta}$ is $\frac{1-\alpha}{\alpha}\cdot\alpha\theta<1$, we have $s\gg 1$. We now split the integration in (43) into three regions. For $u\in(0,s]$, by the local large deviation (31), Karamata’s theorem, and $j(t-u)\leq C/l(t)$,
$\int_{0}^{s}{\mathbb{P}}(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\int_{0}^{s}\frac{CuL\big(B_{t}/\ell(t)^{\theta}\big)}{B_{u}B_{t}^{\alpha}/\ell(t)^{\alpha\theta}}\cdot\frac{1}{l(t)}\mathrm{d}u\leq\frac{Cs^{2}\ell(t)^{\alpha\theta+\varepsilon}}{tB_{s}\hskip 1.0ptl(t)}\leq\frac{C\hskip 1.0pt(t/s)^{\frac{1}{\alpha}-2+\varepsilon}}{\ell(t)^{1-\alpha\theta-\varepsilon}}.$
For $u\in(s,t-s]$, we bound $j(t-u)\leq j(s)\leq C/l(s)$, and use (46) to obtain
$\int^{t-s}_{s}{\mathbb{P}}(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\frac{C}{l(s)}\big(l(t)-l(s)\big)\leq\frac{C}{\ell(t)}(t/s)^{\frac{1}{\alpha}-1+\varepsilon}.$
@@
$\int_{t-s}^{t}{\mathbb{P}}(X^{x}_{u}=0)j(t-u)\mathrm{d}u\leq\frac{C}{B_{t}}\int_{0}^{s}j(u)\mathrm{d}u\leq\frac{Cs}{B_{t}l(s)}\leq\frac{C}{\ell(t)}(t/s)^{-1+\varepsilon}.$
Recalling that $B_{t}$ is $1/\alpha$-varying and $\ell(t)$ is $(1/\alpha-1)$-varying, we have $1/B_{t}\leq C/\ell(t)$. Thus,
${\mathbb{P}}(X^{x}_{t}=0)\leq C/B_{t}\leq C/\ell(t).$
Combining the four bounds above and inserting $t/s=\ell(t)^{\alpha\theta}$ gives